OpenStudy (anonymous):

WIN MEDAL!! A new athletic shoe store had expenses of $40,000 for designing and building the shelves and counters and $120,000 for the first year’s shoe inventory. So far this year, the shoe sales are $70,000. What do the sales have to be for the rest of the year for the store to break even?

OpenStudy (anonymous):

@Miracrown @eric_d !!!

OpenStudy (eric_d):

40k+120k= x 70k-x= ...

OpenStudy (eric_d):

sorry x- 70000= final answer

OpenStudy (anonymous):

I don't get it...

OpenStudy (eric_d):

40000+120000=

OpenStudy (eric_d):

That wld be their total expenses

OpenStudy (anonymous):